OpenStudy (anonymous):

find the area of the region between y=x^1/2 and y=x^1/3 for 0 less than equal to x less than equal to 1.

myininaya (myininaya):

first find the intersections of y=x^(1/2) and y=x^(1/3)

myininaya (myininaya):

by setting x^(1/2)=x^(1/3) solve for x

OpenStudy (anonymous):

so i get 1 and 0?

myininaya (myininaya):

right so now we need to figure out which curve is above the other curve between 0 and 1

myininaya (myininaya):

for example which is true: \[(\frac{1}{2})^\frac{1}{2}> (\frac{1}{2})^\frac{1}{3}? \\ \text{ or is } \\ (\frac{1}{2})^\frac{1}{3} >(\frac{1}{2})^\frac{1}{2}?\]

OpenStudy (anonymous):

i graphed it in my calculator and it appears x^1/2 is above the other

myininaya (myininaya):

i actually think (1/2)^(1/3) is bigger than (1/2)^(1/2)

myininaya (myininaya):

so on that interval from x=0 to x=1 x^(1/3)>x^(1/2)

OpenStudy (anonymous):

ok i got that far and i have it set up as you showed but where do i go from there?

myininaya (myininaya):

\[\int\limits_{0}^{1}(x^\frac{1}{3}-x^\frac{1}{2}) dx\]

myininaya (myininaya):

not on that interval

OpenStudy (anonymous):

My mistake you're right let me delete that!

OpenStudy (anonymous):

oh yea your right because the interval is from 0 to 1

OpenStudy (anonymous):

so were looking at that interval on the graph

OpenStudy (anonymous):

how do i solve now?

myininaya (myininaya):

but if you did choose the first way you could always take the | | of the answer

myininaya (myininaya):

use the integration rule called power rule

OpenStudy (anonymous):

so i get square root of x and cubed root of x?

myininaya (myininaya):

\[\int\limits_{}^{}x^n dx=\frac{x^{n+1}}{n+1} +C , n \neq -1 \]

OpenStudy (anonymous):

so e than have x^1.5/1.5?

OpenStudy (anonymous):

and x^1.3/1.3

OpenStudy (anonymous):

and i solve using the interval of 1 and 0 and plu in to those

OpenStudy (anonymous):

plug*

OpenStudy (anonymous):

is it -1/12

myininaya (myininaya):

\[\int\limits\limits_{0}^{1}(x^\frac{1}{3}-x^\frac{1}{2}) dx \\ =[\frac{x^{\frac{1}{3}+1}}{\frac{1}{3}+1}-\frac{x^{\frac{1}{2}+1}}{\frac{1}{2}+1}]_{0}^{1} \\ =[\frac{x^{\frac{1}{3}+\frac{3}{3}}}{\frac{1}{3}+\frac{3}{3}}-\frac{x^{\frac{1}{2}+\frac{2}{2}}}{\frac{1}{2}+\frac{2}{2}} ]_0^1 \\=[\frac{x^\frac{4}{3}}{\frac{4}{3}} -\frac{x^\frac{3}{2}}{\frac{3}{2}}]^1_0 \\ =[\frac{3}{4}x^\frac{4}{3}-\frac{2}{3}x^\frac{3}{2}]_0^1\] and plug in the limits

myininaya (myininaya):

hmm you should got a positive answer

myininaya (myininaya):

did you do x^(1/2)-x^(1/3)?

myininaya (myininaya):

we are suppose to do greater-smaller

OpenStudy (anonymous):

yea i had them switched whoops!

myininaya (myininaya):

ok well you can just take the | | of the result

OpenStudy (anonymous):

so then its just 1/12?

myininaya (myininaya):

yes 3/4-2/3 is 1/2

myininaya (myininaya):

1/12*

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

what other integration rules are there like the power rule you showed me?

OpenStudy (anonymous):

thank you very much for your help! Have a great day :)

OpenStudy (anonymous):

could you help me with one more?

OpenStudy (anonymous):

find the area of the region between y=x^2 and y=x^3 for 0 less than equal to x less than equal to 1.